find the pressure of 16g of O2 gas in atm present in a flask 1.12 lit at 273K

Grade:11

1 Answers

ADRIAN
7 years ago
Apply pv=nrt .put the given values put R as 0.0821lit atm /k mol. And then solve it mathematically .Pv = nrt P x 1.12 = 1/2 x0.0821 x 273P = 0.0821 x 273 /2.24P = 22.4/2.24P = 10.